The mean free path of nitrogen at 0℃ and 1 atm . is 0.80×10^-5 cm . At this temp. and pressure their are 2.7×10^19 molecules/cm^3 . Calculate the molecule diameter.

Grade:12th pass

2 Answers

Eshan
5 years ago
Dear student,

Use the relation\lambda=\dfrac{RT}{\sqrt{2}\pi d^2N_AP}to obtain the value ofd, whereN_Ais avogadro number.

Regards.
amardeep singh
5 years ago
use the formula (lambda=RT/root2 pi sigma square P Na.    solve by using this formula you should  get answer